OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)? nswer as a function beginning with f(x). Format using this sample answer:f(x)=(3)(x^2-21x-19)

OpenStudy (saifoo.khan):

Since there are 2 roots. -3 and 1 You can write that as: x = -3 and x = 1 x + 3 = 0 and x -1 = 0

OpenStudy (saifoo.khan):

Now you have (x+3)(x-1) = 0 Solve this using foil.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats foil?

OpenStudy (anonymous):

im really bad at math especially this so can you explain please?

OpenStudy (anonymous):

is that the answer? the x^2+2x-3 thing because i need it in the f(x) formant.

OpenStudy (anonymous):

format

OpenStudy (anonymous):

like this f(x)=(3)(x^2-21x-19)

OpenStudy (saifoo.khan):

Umm, yes. f(x) = x^2 +2x -3

OpenStudy (anonymous):

ok thank you tons!